# Decision making in stochastic extensive form I: Stochastic decision forests

### v1

@@ A.2 Section 2
Hence, put aside equalities, $\geq_{\bm{\mathsf{X}}}$ is given by ${\bm{\mathsf{x}}}_{0}\geq_{\bm{\mathsf{X}}}{\bm{\mathsf{x}}}_{k}$, for $k=1,2$. Moreover, the roots of $F$ are given by the image of ${\bm{\mathsf{x}}}_{0}$. It follows that ${\bm{\mathsf{X}}}$ satisfies axioms 3c and 3d.
As the domains of all ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$ are equal to the sample space $\Omega$, the maximality axiom 3e is satisfied. As ${\bm{\mathsf{X}}}$ is finite, the separability condition 3f is fulfilled. ∎
@@
For all $k=0,1$, ${\bm{\mathsf{x}}}_{k}$ is defined on $\Omega$ as in the previous lemma. For $k=2$, it has to be noted that $\{\omega_{2}\}\in\mathscr{A}$. Clearly, ${\bm{\mathsf{x}}}^{\prime}$ is a section of $\pi$ for all ${\bm{\mathsf{x}}}^{\prime}\in{\bm{\mathsf{X}}}^{\prime}$. The union of the images of the elements of ${\bm{\mathsf{X}}}^{\prime}$ is the set of nodes with at least two elements. Again, as $\{\{w^{\prime}\}\mid w^{\prime}\in W\}\subseteq F^{\prime}$, this union equals $X^{\prime}$. It follows that ${\bm{\mathsf{X}}}^{\prime}$ satisfies axioms 3a and 3b.
@@
Hence, put aside equalities, $\geq_{{\bm{\mathsf{X}}}^{\prime}}$ is given by ${\bm{\mathsf{x}}}^{\prime}_{0}\geq_{{\bm{\mathsf{X}}}^{\prime}}{\bm{\mathsf{x}}}^{\prime}_{1}$ and ${\bm{\mathsf{x}}}^{\prime}_{0}\geq_{{\bm{\mathsf{X}}}^{\prime}}{\bm{\mathsf{x}}}^{\prime}_{2}$. The roots of $F^{\prime}$ are given by the image of ${\bm{\mathsf{x}}}^{\prime}_{0}$. It follows that ${\bm{\mathsf{X}}}^{\prime}$ satisfied axioms 3c and 3d.
The maximality axiom 3e requires only a little bit more justification this time. If $\bar{\bm{\mathsf{X}}}^{\prime}$ satisfies the first four axioms 3a, 3b, 3c, and 3d, and ${\bm{\mathsf{X}}}^{\prime}$ refines $\bar{\bm{\mathsf{X}}}^{\prime}$, then ${\bm{\mathsf{x}}}^{\prime}_{0},{\bm{\mathsf{x}}}^{\prime}_{1}\in\bar{\bm{\mathsf{X}}}^{\prime}$ because of the full domain of these presumed random moves and axioms 3a and 3b. By proposition 2.11 and remark 2.13, there there must be one and only one other element of $\bar{\bm{\mathsf{X}}}^{\prime}$, namely ${\bm{\mathsf{x}}}^{\prime}_{2}$. Hence, $\bar{\bm{\mathsf{X}}}^{\prime}={\bm{\mathsf{X}}}^{\prime}$. The separability condition 3f is again fulfilled because of the finiteness of the situation. ∎
@@
(Ad $\Rightarrow$”):  Suppose AP.W1 to hold and let $x\in F$ be not a singleton. Then, there are $(t,w)\in\mathbb{T}\times W$ with $x=x_{t}(w)$. Let $(t^{\prime},w^{\prime})\in\mathbb{T}\times W$ such that $x=x_{t^{\prime}}(w^{\prime})$. Hence, $w\in x_{t^{\prime}}(w^{\prime})$ and thus $x=x_{t^{\prime}}(w)$. By assumption AP.W1, $t=t^{\prime}$. Hence, $\mathbb{T}_{x}=\{t\}$.
@@
(Ad axiom 1):  As $F\subseteq\mathcal{P}(W)$, $F$ is a $W$-poset. We show that it is a “bounded” and “irreducible” “$W$-set tree”, in the language of [3], and that any principal up-set in $(F,\supseteq)$ has a maximal element. From this, using [3, theorem 3], we obtain, expressed in the language of the present paper, that $F$ is decision forest over $W$.
@@
Regarding “irreducibility”, let $w_{0},w_{1}\in W$ with $w_{0}\neq w_{1}$. Then $\{w_{k}\}\in F$ satisfy $w_{k}\in\{w_{k}\}\setminus\{w_{1-k}\}$ for both $k=0,1$. Whence irreducibility.
@@
Let $\sup\mathbb{T}^{\prime}$ denote the supremum of $\mathbb{T}^{\prime}$ in ¯$\bar{\mathbb{R}}_{+}=[0,+\infty]$. As $c$ is a chain and by definition of $\mathbb{T}^{\prime}$, for all $t\in\mathbb{T}$ with $t<\sup\mathbb{T}^{\prime}$, there is a unique $a(t)\in\mathbb{A}$ such that for all $t^{\prime}\in\mathbb{T}^{\prime}$ with $t<t^{\prime}$ and $f_{t^{\prime}}\in\mathbb{A}^{\mathbb{T}}$ satisfying $x_{t^{\prime}}(\omega,f_{t^{\prime}})\in c$ we have $f_{t^{\prime}}(t)=a(t)$. Let $\tilde{f}\in\mathbb{A}^{\mathbb{T}}$ be such that $\tilde{f}(t)=a(t)$ for all $t\in[0,\sup\mathbb{T}^{\prime})\cap\mathbb{T}$. Hence, all $t^{\prime}\in\mathbb{T}^{\prime}$ satisfy $x_{t^{\prime}}(\omega,\tilde{f})\in c\subseteq F$. By assumption AP.W2, there is $f\in\mathbb{A}^{\mathbb{T}}$ such that $(\omega,f)\in W$ and $x_{t^{\prime}}(\omega,f)=x_{t^{\prime}}(\omega,\tilde{f})\in c$ for all $t^{\prime}\in\mathbb{T}^{\prime}$.
@@
(Ad axiom 3a):  By definition of the sets $D_{t,f}$, $(t,f)\in\mathbb{T}\times\mathbb{A}^{\mathbb{T}}$, and in view of the fact that $F$ contains all singletons in $\mathcal{P}(W)$, all elements of ${\bm{\mathsf{X}}}$ map into the set $X$ of moves of the decision forest $F$ over $W$. Moreover, by assumption AP.W0, $D_{t,f}\in\mathscr{A}$. Furthermore, any ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$ is a section of $\pi$. For this represent ${\bm{\mathsf{x}}}$ as ${\bm{\mathsf{x}}}_{t}(f)$ for $(t,f)\in\mathbb{T}\times\mathbb{A}^{\mathbb{T}}$ such that $D_{t,f}\neq\emptyset$, and let $\omega\in D_{t,f}$. Then,
@@
(Ad axiom 3b):  Let $x\in X$. Then $x$ has at least two elements, because $F$ is a decision forest over $W$. Hence, there are $(\omega,f)\in W$ and $t\in\mathbb{T}$ such that $x=x_{t}(\omega,f)$. Hence, $\omega\in D_{t,f}$, and $x={\bm{\mathsf{x}}}_{t}(f)(\omega)$.
(Ad axiom 3c):  Let ${\bm{\mathsf{x}}}_{1},{\bm{\mathsf{x}}}_{2}\in{\bm{\mathsf{X}}}$ such that there is $\omega\in D_{{\bm{\mathsf{x}}}_{1}}\cap D_{{\bm{\mathsf{x}}}_{2}}$ with ${\bm{\mathsf{x}}}_{1}(\omega)\supseteq{\bm{\mathsf{x}}}_{2}(\omega)$. Hence, there are $t_{1},t_{2}\in\mathbb{T}$ and $f\in\mathbb{A}^{\mathbb{T}}$ such that $\omega\in D_{t_{k},f}$ and ${\bm{\mathsf{x}}}_{k}={\bm{\mathsf{x}}}_{t_{k}}(\omega,f)$, for both $k=1,2$. If we had $t_{1}>t_{2}$, then, by construction of the nodes and by assumption AP.W1, ${\bm{\mathsf{x}}}_{1}(\omega)=x_{t_{1}}(\omega,f)\subsetneq x_{t_{2}}(\omega,f)={\bm{\mathsf{x}}}_{2}(\omega)$, because they contain at least two elements as moves – a contradiction. Hence, $t_{1}\leq t_{2}$.
@@
(Ad axiom 3d):  Let ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$ admitting $\omega\in D_{\bm{\mathsf{x}}}$ with ${\bm{\mathsf{x}}}(\omega)=(\{\omega\}\times\mathbb{A}^{\mathbb{T}})\cap W$. Hence, for any $f\in\mathbb{A}^{\mathbb{T}}$, $D_{0,f}\neq\emptyset$ and ${\bm{\mathsf{x}}}_{0}(f)(\omega)={\bm{\mathsf{x}}}(\omega)$. Let us pick an $f\in\mathbb{A}^{\mathbb{T}}$. Hence, ${\bm{\mathsf{x}}}_{0}(f)\in{\bm{\mathsf{X}}}$ and, by axiom 3c, ${\bm{\mathsf{x}}}={\bm{\mathsf{x}}}_{0}(f)$. Thus, for all $\omega^{\prime}\in D_{\bm{\mathsf{x}}}=D_{0,f}$, we have ${\bm{\mathsf{x}}}(\omega^{\prime})=x_{0}(\omega^{\prime},f)=(\{\omega^{\prime}\}\times\mathbb{A}^{\mathbb{T}})\cap W$.
(Ad axiom 3e):  Let $\bar{\bm{\mathsf{X}}}$ be a set satisfying the first four axioms of 3 that is refined by ${\bm{\mathsf{X}}}$. By definition of the latter, for any $\bar{\bm{\mathsf{x}}}\in\bar{\bm{\mathsf{X}}}$ there is $P_{\bar{\bm{\mathsf{x}}}}\subseteq{\bm{\mathsf{X}}}$ such that $\bar{\bm{\mathsf{x}}}=\bigcup P_{\bar{\bm{\mathsf{x}}}}$. As $\bar{\bm{\mathsf{X}}}$ and ${\bm{\mathsf{X}}}$ induce partitions of $X$, by proposition 2.11 and remark 2.13, and as $D_{\bm{\mathsf{x}}}\neq\emptyset$ for all ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$ by construction of ${\bm{\mathsf{X}}}$, we infer that there is a unique map $b\colon{\bm{\mathsf{X}}}\to\bar{\bm{\mathsf{X}}},{\bm{\mathsf{x}}}\mapsto b({\bm{\mathsf{x}}})$ such that $D_{\bm{\mathsf{x}}}\subseteq D_{b({\bm{\mathsf{x}}})}$ and $b({\bm{\mathsf{x}}})|_{D_{\bm{\mathsf{x}}}}={\bm{\mathsf{x}}}$, for all ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$. For each pair $(t,f)\in\mathbb{T}\times\mathbb{A}^{\mathbb{T}}$ with $D_{t,f}\neq\emptyset$, denote $\bar{\bm{\mathsf{x}}}_{t}(f)=b({{\bm{\mathsf{x}}}_{t}(f)})$ the image of ${\bm{\mathsf{x}}}_{t}(f)$ under this map.
@@
Indeed, otherwise there would be $\omega\in D_{t,f}\cap D_{u,g}$ for that that definition of $\bar{x}_{t}(f)$ combined with ()($\bullet$)( ∙ ) would imply $x_{t}(\omega,f)=x_{u}(\omega,g)$. Hence, for $v=t\wedge u$, we would have $f|_{[0,v)_{\mathbb{T}}}=g|_{[0,v)_{\mathbb{T}}}$ and there would be $w\in W$ with $x_{t}(\omega,f)=x_{t}(w)=x_{u}(w)=x_{u}(\omega,g)$. As these are moves, assumption AP.W1 would imply $t=u=v$, whence ${\bm{\mathsf{x}}}_{t}(f)={\bm{\mathsf{x}}}_{u}(g)$ which is absurd. Hence, we would necessarily have $D_{t,f}\cap D_{u,g}=\emptyset$.
We can assume, without loss of generality, that $t\leq u$. There would be two cases. First, $D_{t,f}\cap D_{t,g}\neq\emptyset$. Then, $f|_{[0,t)_{\mathbb{T}}}\neq g|_{[0,t)_{\mathbb{T}}}$, because else $D_{t,f}=D_{t,g}$ whence $D_{u,g}\subseteq D_{t,g}=D_{t,f}$ which contradicts ()($\dagger$)( † ). The second case would be $D_{t,f}\cap D_{t,g}=\emptyset$. Then, by assumption AP.W3, there would be $v\in[0,t)_{\mathbb{T}}$ such that
@@
Let $\omega^{\prime}\in D_{v,f}\cap D_{v,g}$. Then, $x_{v}(\omega^{\prime},f)$ and $x_{v}(\omega^{\prime},g)$ would not be comparable in $(F,\supseteq)$ since equality would imply $f|_{[0,v)_{\mathbb{T}}}=g|_{[0,v)_{\mathbb{T}}}$. Hence, by the definition of $\geq_{\bar{\bm{\mathsf{X}}}}$, the same would be true for $\bar{\bm{\mathsf{x}}}_{v}(f)$ and $\bar{\bm{\mathsf{x}}}_{v}(g)$ in $(\bar{\bm{\mathsf{X}}},\geq_{\bar{\bm{\mathsf{X}}}})$. But, on the other hand, ${\bm{\mathsf{x}}}_{v}(f)\geq_{\bm{\mathsf{X}}}{\bm{\mathsf{x}}}_{t}(f)$ and ${\bm{\mathsf{x}}}_{v}(g)\geq_{\bm{\mathsf{X}}}{\bm{\mathsf{x}}}_{u}(g)$, whence by definition of $\geq_{\bm{\mathsf{X}}}$ and axiom 3c applied to $\bar{\bm{\mathsf{X}}}$, we get $\bar{\bm{\mathsf{x}}}_{v}(f),\bar{\bm{\mathsf{x}}}_{v}(g)\in\uparrow\bar{\bm{\mathsf{x}}}_{t}(f)$. As $(\bar{\bm{\mathsf{X}}},\geq_{\bar{\bm{\mathsf{X}}}})$ is a tree by proposition 2.11 and remark 2.13, $\bar{\bm{\mathsf{x}}}_{v}(f)$ and $\bar{\bm{\mathsf{x}}}_{v}(g)$ would have to be comparable, a contradiction. Hence, the initial assumption was false, and we infer that $P_{\bar{\bm{\mathsf{x}}}}=\{\bar{\bm{\mathsf{x}}}\}$ for all $\bar{\bm{\mathsf{x}}}\in\bar{\bm{\mathsf{X}}}$.
@@
(Ad axiom 3f):  As $\mathbb{T}$ is embedded in $\mathbb{R}_{+}$, there is countable $\mathbb{T}_{0}\subseteq\mathbb{T}$, dense in $\mathbb{T}$ with respect to the induced topology on $\mathbb{T}$. Let ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$. Then, by construction of the nodes there are $t\in\mathbb{T}$ and $f\in\mathbb{A}^{\mathbb{T}}$ such that ${\bm{\mathsf{x}}}={\bm{\mathsf{x}}}_{t}(f)$. Let
${\bm{\mathsf{X}}}_{0}=\{{\bm{\mathsf{x}}}_{t_{0}}(f)\mid t_{0}\in\mathbb{T}_{0},\leavevmode\nobreak\ t_{0}<t\}.$
Let ${\bm{\mathsf{x}}}^{\prime}\in{\bm{\mathsf{X}}}$ such that ${\bm{\mathsf{x}}}^{\prime}<_{\bm{\mathsf{X}}}{\bm{\mathsf{x}}}$. Then, by construction of the nodes and lemma 2.21, there is $t^{\prime}\in\mathbb{T}$ with $t^{\prime}<t$ such that ${\bm{\mathsf{x}}}^{\prime}={\bm{\mathsf{x}}}_{t^{\prime}}(f)$. Hence, there is $t_{0}\in\mathbb{T}_{0}$ with $t^{\prime}\leq t_{0}\leq t$, and we easily obtain, directly applying the definitions, $D_{t^{\prime},f}\supseteq D_{t_{0},f}\supseteq D_{t,f}$ and for all $\omega\in D_{t,f}$, $x_{t_{0}}(\omega,f)\supseteq x_{t}(\omega,f)$, and for all $\omega_{0}\in D_{t_{0},f}$, $x_{t^{\prime}}(\omega_{0},f)\supseteq x_{t_{0}}(\omega_{0},f)$. Thus $D_{t_{0},f}\neq\emptyset$, ${\bm{\mathsf{x}}}_{t_{0}}(f)\in{\bm{\mathsf{X}}}$, and ${\bm{\mathsf{x}}}^{\prime}\geq_{\bm{\mathsf{X}}}{\bm{\mathsf{x}}}_{t_{0}}(f)\geq_{\bm{\mathsf{X}}}{\bm{\mathsf{x}}}$. ∎

### v2

@@ A.2 Section 2
Hence, put aside equalities, $\geq_{\bm{\mathsf{X}}}$ is given by ${\bm{\mathsf{x}}}_{0}\geq_{\bm{\mathsf{X}}}{\bm{\mathsf{x}}}_{k}$, for $k=1,2$. Moreover, the roots of $F$ are given by the image of ${\bm{\mathsf{x}}}_{0}$. It follows that ${\bm{\mathsf{X}}}$ satisfies Axioms 4(a) and 4(b).
As the domains of all ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$ are equal to the sample space $\Omega$, the maximality property of Axiom 4(c) is satisfied. ∎
@@
For all $k=0,1$, ${\bm{\mathsf{x}}}_{k}$ is defined on $\Omega$ as in the previous lemma. For $k=2$, it has to be noted that $\{\omega_{2}\}\in\mathscr{E}$. Clearly, ${\bm{\mathsf{x}}}^{\prime}$ is a section of $\pi$ for all ${\bm{\mathsf{x}}}^{\prime}\in{\bm{\mathsf{X}}}^{\prime}$. The union of the images of the elements of ${\bm{\mathsf{X}}}^{\prime}$ is the set of nodes with at least two elements. Again, as $\{\{w^{\prime}\}\mid w^{\prime}\in W\}\subseteq F^{\prime}$, this union equals $X^{\prime}$. It follows that ${\bm{\mathsf{X}}}^{\prime}$ satisfies Axioms 33(a) and 33(b).
@@
Hence, put aside equalities, $\geq_{{\bm{\mathsf{X}}}^{\prime}}$ is given by ${\bm{\mathsf{x}}}^{\prime}_{0}\geq_{{\bm{\mathsf{X}}}^{\prime}}{\bm{\mathsf{x}}}^{\prime}_{1}$ and ${\bm{\mathsf{x}}}^{\prime}_{0}\geq_{{\bm{\mathsf{X}}}^{\prime}}{\bm{\mathsf{x}}}^{\prime}_{2}$. The roots of $F^{\prime}$ are given by the image of ${\bm{\mathsf{x}}}^{\prime}_{0}$. It follows that ${\bm{\mathsf{X}}}^{\prime}$ satisfied Axioms 4(a) and 4(b).
The maximality property of Axiom 4(c) requires only a little bit more justification this time. If $\bar{\bm{\mathsf{X}}}^{\prime}$ is such that $(F,\pi,\bar{\bm{\mathsf{X}}}^{\prime})$ is an sdf and ${\bm{\mathsf{X}}}^{\prime}$ refines $\bar{\bm{\mathsf{X}}}^{\prime}$, then any $\bar{\bm{\mathsf{x}}}^{\prime}\in\bar{\bm{\mathsf{X}}}^{\prime}$ with values in $\operatorname{im}{\bm{\mathsf{x}}}^{\prime}_{1}\cup\operatorname{im}{\bm{\mathsf{x}}}^{\prime}_{0}$, we have $\bar{\bm{\mathsf{x}}}^{\prime}\in\{{\bm{\mathsf{x}}}^{\prime}_{1},{\bm{\mathsf{x}}}^{\prime}_{0}\}$ because the domain of ${\bm{\mathsf{x}}}^{\prime}_{1}$ and ${\bm{\mathsf{x}}}^{\prime}_{0}$ is $\Omega$. Hence, there can and must be exactly one other element of $\bar{\bm{\mathsf{X}}}^{\prime}$, namely ${\bm{\mathsf{x}}}^{\prime}_{2}$. Hence, $\bar{\bm{\mathsf{X}}}^{\prime}={\bm{\mathsf{X}}}^{\prime}$. ∎
@@
$(F,\pi,{\bm{\mathsf{X}}})$ is easily shown to be an sdf, similarly as above.
$\rho$ is surjective by assumption. Let $\omega_{k}\in\rho^{-1}(\{k\})$, for $k=1,2$. Then for both of these $k$, ${\bm{\mathsf{x}}}_{k}(\omega_{k})\supsetneq{\bm{\mathsf{x}}}_{3-k}(\omega_{k})$, but, as a consequence, ${\bm{\mathsf{x}}}_{k}\ngeq_{\bm{\mathsf{X}}}{\bm{\mathsf{x}}}_{3-k}$. Hence, $(F,\pi,{\bm{\mathsf{X}}})$ is not order consistent. ∎
Proof of Lemma 2.17.
(Ad $\Rightarrow$”):  Suppose AP.SDF1 to hold and let $x\in F$ be not a singleton. Then, there are $(t,w)\in\mathbb{T}\times W$ with $x=x_{t}(w)$. Let $(t^{\prime},w^{\prime})\in\mathbb{T}\times W$ such that $x=x_{t^{\prime}}(w^{\prime})$. Hence, $w\in x_{t^{\prime}}(w^{\prime})$ and thus $x=x_{t^{\prime}}(w)$. By Assumption AP.SDF1, $t=t^{\prime}$. Hence, $\mathbb{T}_{x}=\{t\}$.
@@
(Ad Axiom 1):  As $F\subseteq\mathcal{P}(W)$, $F$ is a $W$-poset. We show that it is a “bounded” and “irreducible” “$W$-set tree”, in the language of [3], and that any principal up-set in $(F,\supseteq)$ has a maximal element. From this, using [3, Theorem 3], we obtain, expressed in the language of the present paper, that $F$ is decision forest on $W$.
@@
Regarding “irreducibility”, let $w_{0},w_{1}\in W$ with $w_{0}\neq w_{1}$. Then $\{w_{k}\}\in F$ satisfies $w_{k}\in\{w_{k}\}\setminus\{w_{1-k}\}$ for both $k=0,1$. Whence irreducibility.
@@
Let $\tilde{\mathbb{T}}=\{t\in\mathbb{T}\mid\exists t^{\prime}\in\mathbb{T}^{\prime}\colon t<t^{\prime}\}$, a convex subset of $\mathbb{T}$. As $c$ is a chain and by definition of $\mathbb{T}^{\prime}$, for all $t\in\tilde{\mathbb{T}}$, there is a unique $a(t)\in\mathbb{A}$ such that for all $t^{\prime}\in\mathbb{T}^{\prime}$ with $t<t^{\prime}$ and $f_{t^{\prime}}\in\mathbb{A}^{\mathbb{T}}$ satisfying $x_{t^{\prime}}(\omega,f_{t^{\prime}})\in c$ we have $f_{t^{\prime}}(t)=a(t)$. Let $\tilde{f}\in\mathbb{A}^{\mathbb{T}}$ be such that $\tilde{f}(t)=a(t)$ for all $t\in\tilde{\mathbb{T}}$. Hence, all $t^{\prime}\in\mathbb{T}^{\prime}$ satisfy $x_{t^{\prime}}(\omega,\tilde{f})\in c\subseteq F$. By Assumption AP.SDF2, there is $f\in\mathbb{A}^{\mathbb{T}}$ such that $(\omega,f)\in W$ and $x_{t^{\prime}}(\omega,f)=x_{t^{\prime}}(\omega,\tilde{f})\in c$ for all $t^{\prime}\in\mathbb{T}^{\prime}$.
@@
(Ad Axiom 33(a)):  By definition of the sets $D_{t,f}$, $(t,f)\in\mathbb{T}\times\mathbb{A}^{\mathbb{T}}$, and in view of the fact that $F$ contains all singletons in $\mathcal{P}(W)$, all elements of ${\bm{\mathsf{X}}}$ map into the set $X$ of moves of the decision forest $F$ on $W$. Moreover, by Assumption AP.SDF0, $D_{t,f}\in\mathscr{E}$. Furthermore, any ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$ is a section of $\pi$. For this represent ${\bm{\mathsf{x}}}$ as ${\bm{\mathsf{x}}}_{t}(f)$ for $(t,f)\in\mathbb{T}\times\mathbb{A}^{\mathbb{T}}$ such that $D_{t,f}\neq\emptyset$, and let $\omega\in D_{t,f}$. Then,
@@
(Ad Axiom 33(b)):  Let $x\in X$. Then $x$ has at least two elements, because $F$ is a decision forest on $W$. Hence, there are $(\omega,f)\in W$ and $t\in\mathbb{T}$ such that $x=x_{t}(\omega,f)$. Hence, $\omega\in D_{t,f}$, and $x={\bm{\mathsf{x}}}_{t}(f)(\omega)$.
(Ad Axiom 4(a)):  Let ${\bm{\mathsf{x}}}_{1},{\bm{\mathsf{x}}}_{2}\in{\bm{\mathsf{X}}}$ such that there is $\omega\in D_{{\bm{\mathsf{x}}}_{1}}\cap D_{{\bm{\mathsf{x}}}_{2}}$ with ${\bm{\mathsf{x}}}_{1}(\omega)\supseteq{\bm{\mathsf{x}}}_{2}(\omega)$. Hence, there are $t_{1},t_{2}\in\mathbb{T}$ and $f\in\mathbb{A}^{\mathbb{T}}$ such that $\omega\in D_{t_{k},f}$ and ${\bm{\mathsf{x}}}_{k}={\bm{\mathsf{x}}}_{t_{k}}(\omega,f)$, for both $k=1,2$. If we had $t_{1}>t_{2}$, then, by construction of the nodes and by Assumption AP.SDF1, ${\bm{\mathsf{x}}}_{1}(\omega)=x_{t_{1}}(\omega,f)\subsetneq x_{t_{2}}(\omega,f)={\bm{\mathsf{x}}}_{2}(\omega)$, because they contain at least two elements as moves – a contradiction. Hence, $t_{1}\leq t_{2}$.
@@
(Ad Axiom 4(c)):  Let $\bar{\bm{\mathsf{X}}}$ be a set such that $(F,\pi,\bar{\bm{\mathsf{X}}})$ is an order consistent sdf and that is refined by ${\bm{\mathsf{X}}}$. By definition of the latter, for any $\bar{\bm{\mathsf{x}}}\in\bar{\bm{\mathsf{X}}}$ there is $P_{\bar{\bm{\mathsf{x}}}}\subseteq{\bm{\mathsf{X}}}$ such that $\bar{\bm{\mathsf{x}}}=\bigcup P_{\bar{\bm{\mathsf{x}}}}$. As ${\bm{\mathsf{X}}}$ and $\bar{\bm{\mathsf{X}}}$ induce partitions of $X$, by Proposition 2.12, and as $D_{\bm{\mathsf{x}}}\neq\emptyset$ for all ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$ by construction of ${\bm{\mathsf{X}}}$, we infer that there is a unique map $b\colon{\bm{\mathsf{X}}}\to\bar{\bm{\mathsf{X}}},{\bm{\mathsf{x}}}\mapsto b({\bm{\mathsf{x}}})$ such that $D_{\bm{\mathsf{x}}}\subseteq D_{b({\bm{\mathsf{x}}})}$ and $b({\bm{\mathsf{x}}})|_{D_{\bm{\mathsf{x}}}}={\bm{\mathsf{x}}}$, for all ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$. For each pair $(t,f)\in\mathbb{T}\times\mathbb{A}^{\mathbb{T}}$ with $D_{t,f}\neq\emptyset$, denote $\bar{\bm{\mathsf{x}}}_{t}(f)=b({{\bm{\mathsf{x}}}_{t}(f)})$ the image of ${\bm{\mathsf{x}}}_{t}(f)$ under this map.
@@
Indeed, otherwise there would be $\omega\in D_{t,f}\cap D_{u,g}$ for that we would obtain $x_{t}(\omega,f)=\bar{\bm{\mathsf{x}}}(\omega)=x_{u}(\omega,g)$. Hence, for $v=t\wedge u$, we would have $f|_{[0,v)_{\mathbb{T}}}=g|_{[0,v)_{\mathbb{T}}}$ and there would be $w\in W$ with $x_{t}(\omega,f)=x_{t}(w)=x_{u}(w)=x_{u}(\omega,g)$. As these are moves, Assumption AP.SDF1 would imply $t=u=v$, whence ${\bm{\mathsf{x}}}_{t}(f)={\bm{\mathsf{x}}}_{u}(g)$ which is absurd. Hence, we would necessarily have $D_{t,f}\cap D_{u,g}=\emptyset$.
We can assume, without loss of generality, that $t\leq u$. There would be two cases. First, $D_{t,f}\cap D_{t,g}\neq\emptyset$. Then, $f|_{[0,t)_{\mathbb{T}}}\neq g|_{[0,t)_{\mathbb{T}}}$, because else $D_{t,f}=D_{t,g}$ whence $D_{u,g}\subseteq D_{t,g}=D_{t,f}$ which contradicts ()($\dagger$)( † ). The second case would be $D_{t,f}\cap D_{t,g}=\emptyset$. Then, by Assumption AP.SDF3, there would be $v\in[0,t)_{\mathbb{T}}$ such that
@@
Let $\omega^{\prime}\in D_{v,f}\cap D_{v,g}$. Then, $x_{v}(\omega^{\prime},f)$ and $x_{v}(\omega^{\prime},g)$ would not be comparable in $(F,\supseteq)$ since this would imply $f|_{[0,v)_{\mathbb{T}}}=g|_{[0,v)_{\mathbb{T}}}$. Hence, by the definition of $\geq_{\bar{\bm{\mathsf{X}}}}$, the same would be true for $\bar{\bm{\mathsf{x}}}_{v}(f)$ and $\bar{\bm{\mathsf{x}}}_{v}(g)$ in $(\bar{\bm{\mathsf{X}}},\geq_{\bar{\bm{\mathsf{X}}}})$. But, on the other hand, ${\bm{\mathsf{x}}}_{v}(f)\geq_{\bm{\mathsf{X}}}{\bm{\mathsf{x}}}_{t}(f)$ and ${\bm{\mathsf{x}}}_{v}(g)\geq_{\bm{\mathsf{X}}}{\bm{\mathsf{x}}}_{u}(g)$, whence by definition of $\geq_{\bm{\mathsf{X}}}$ and Axiom 4(a) applied to $\bar{\bm{\mathsf{X}}}$, we get $\bar{\bm{\mathsf{x}}}_{v}(f),\bar{\bm{\mathsf{x}}}_{v}(g)\in\uparrow\bar{\bm{\mathsf{x}}}_{t}(f)$. As $(\bar{\bm{\mathsf{X}}},\geq_{\bar{\bm{\mathsf{X}}}})$ is a forest (which follows from Proposition 2.12 as in the proof of Theorem 2.13), $\bar{\bm{\mathsf{x}}}_{v}(f)$ and $\bar{\bm{\mathsf{x}}}_{v}(g)$ would have to be comparable, a contradiction. Hence, the initial assumption was false, and we infer that $P_{\bar{\bm{\mathsf{x}}}}=\{\bar{\bm{\mathsf{x}}}\}$ for all $\bar{\bm{\mathsf{x}}}\in\bar{\bm{\mathsf{X}}}$.
@@
(Ad statement about sure non-triviality):  For all $w\in W$, we have $x_{0}(w)\supseteq\{w\}$ and $\{w\}\in F$. Hence, for all $\omega\in\Omega$ and $f\in\mathbb{A}^{\mathbb{T}}$, $x_{0}(\omega,f)\in X$ iff $\omega\in D_{0,f}$. It is moreover obvious from the definitions that all $\omega\in\Omega$ and $f\in\mathbb{A}^{\mathbb{T}}$ satisfy $x_{0}(\omega,f)=W_{\omega}$. Hence, for any $f\in\mathbb{A}^{\mathbb{T}}$,
$\{\omega\in\Omega\mid W_{\omega}\in X\}=D_{0,f}=\big{\{}\omega\in\Omega\mid\exists f^{\prime},f^{\prime\prime}\in\mathbb{A}^{\mathbb{T}}\colon[f^{\prime}\neq f^{\prime\prime}\text{ and }(\omega,f^{\prime}),(\omega,f^{\prime\prime})\in W]\big{\}},$
from which the claimed equivalence directly follows in view of ()($\ast$)( ∗ ). ∎